40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2630     Entry time: Tue Feb 23 06:47:57 2010
Author: Koji 
Type: Update 
Category: General 
Subject: IFO situations / low power MC lock 

Work on 22nd Monday:

[MC recovery]

- Tried to lock MC after the computer recovery by Joe.
- A lot of higher modes. I can touch the input periscope or the MC mirrors.
- First tried to align the MC mirrors. MC1 was aligned against the MC REFL PD. MC2/3 was aligned to maximize the transmitted power.
- After the alignment, I got the MC Trans Sum ~8V. Also I saw the flashing of the arm cavities. I decided to take this alignment although the beam looks little bit clipped by the faraday.

[IFO alignment recovery]

- Aligned the arms for TEM00 manually.
- Arm alignment script seems not working now. This could come from the move of the end QPDs
- PRMI/DRMI were aligned. All alignment values saved.

[Low power MC]

[Optical config]

- I fixed the MCT CCD camera. It is quite useful to align the MC.

- Inserted HWP+Cube PBS+HWP combo in the MC incident path.
- First HWP and PBS adjust the light power. The second HWP is fixed at 342deg such that it restores the poralization to S.
- The incident power was measured by the SCIENTECH power meter. Offset of 3mW was subtracted in the table below.

HWP1 angle P_MC_incident comment
126deg 1.03W Max
100 0.39  
90 0.098 Low power max
85 0.021 Low power nominal

- HWP1 85deg is the nominal.

- I needed to touch the steering mirror (indicated by the picture) to obtain TEM00.
  The alignment of the HWPs and the cube PBS didn't change the mode. Thermal lense of the cube?

- I could not lock the MC with the incident power below 100mW. So the BS in the MC REFL path was replaced by a total reflector (Y1-45S).
- This increased the power on the MC REFL PD x10 of the previous. NOW WE ARE CONSTRAINED BETWEEN 81deg~90deg. DON'T ROTATE FURTHER!
- The original BS was stored on the AP table as shown in the picture.
- This total reflector disabled the MC WFS QPDs. We can't use them.

[Lock of the MC with 20mW incident]
- Disable the MC autolocker.
- Disable the MC WFS.
- Run
  /cvs/cds/caltech/scripts/MC/mcloopson
- Turn on the MCL servo.
- Set the MCL gain to 1.5 (it was nominally 0.3 for the high power)
- Just wait until lock.

[Gain boost after the lock] ...If you like to have more gain
- There was almost no room to increase the MCL gain.
- MC_REFL_GAIN can be increased from +6dB to +20dB
  ezcawrite "C1:IOO-MC_REFL_GAIN" 20
- MC_VCO_GAIN can be increased from -3dB to +2dB
  ezcawrite "C1:IOO-MC_VCO_GAIN" 2
- Crank the FSS gains
  ezcawrite C1:PSL-FSS_MGAIN `ezcaread -n C1:PSL-STAT_FSS_NOM_C_GAIN`
  ezcawrite C1:PSL-FSS_FASTGAIN `ezcaread -n C1:PSL-STAT_FSS_NOM_F_GAIN`

[If lock is lost]
- Run
  /cvs/cds/caltech/scripts/MC/mcdown

Attachment 1: MC_incident.png  467 kB  Uploaded Tue Feb 23 12:29:21 2010  | Show | Show all
Attachment 2: MC_REFL.png  519 kB  Uploaded Tue Feb 23 12:29:30 2010  | Show | Show all
ELOG V3.1.3-